OpenStudy (anonymous):

Once a crate is in motion, a horizontal force of 53 N keeps the crate moving with a constant velocity. Find the coefficient of kinetic friction between the crate and the floor.

OpenStudy (anonymous):

\[F _{f} = \mu N\] \[N = mg\] \[\mu = \frac{ F_{f} }{ mg }\] Given \[g = 9.8 m/s^2\] you would get \[\mu = \frac{ 53 }{ m*9.8 }\] You're missing the mass of the crate(m).

OpenStudy (anonymous):

thank you, i realized the mass was on a different page lol